# Multi-LF: A Unified Continuous Learning Framework for Real-Time DDoS Detection in Multi-Environment Networks

### v1

@@ Abstract
Detecting Distributed Denial of Service (DDoS) attacks in Multi-Environment (M-En) networks presents significant challenges due to diverse malicious traffic patterns and the evolving nature of cyber threats. Existing AI-based detection systems struggle to adapt to new attack strategies and lack real-time attack detection capabilities with high accuracy and efficiency. This study proposes an online, continuous learning methodology for DDoS detection in M-En networks, enabling continuous model updates and real-time adaptation to emerging threats, including zero-day attacks. First, we develop a unique M-En network dataset by setting up a realistic, real-time simulation using the NS-3 tool, incorporating both victim and bot devices. DDoS attacks with varying packet sizes are simulated using the DDoSim application across IoT and traditional IP-based environments under M-En network criteria. Our approach employs a multi-level framework (Multi-LF) featuring two machine learning models: a lightweight Model 1 (M1) trained on a selective, critical packet dataset for fast and efficient initial detection, and a more complex, highly accurate Model 2 (M2) trained on extensive data. When M1 exhibits low confidence in its predictions, the decision is escalated to M2 for verification and potential fine-tuning of M1 using insights from M2. If both models demonstrate low confidence, the system flags the incident for human intervention, facilitating model updates with human-verified categories to enhance adaptability to unseen attack patterns. We validate the Multi-LF through real-world simulations, demonstrating superior classification accuracy of 0.999 and low prediction latency of 0.866 seconds compared to established baselines. Furthermore, we evaluate performance in terms of memory usage (3.632 MB) and CPU utilization (10.05%) in real-time scenarios.
Index Terms:
@@
DDoS Attacks, Network Emulation, Dataset Collection, Continuous Learning, M-En Networks, Zero-Day Attacks
I Introduction
Distributed Denial-of-Service (DDoS) attacks pose a critical challenge to network security by overwhelming target servers or networks with excessive traffic, thereby depleting their bandwidth, memory, and processing resources [1]. Attackers employ a variety of techniques to orchestrate these attacks. One prevalent method involves the deployment of botnets—networks of compromised devices such as Internet of Things (IoT) devices infected with malware—which are remotely controlled to generate vast amounts of traffic directed at the target [2]. In addition, adversaries exploit inherent properties of certain network protocols by involving intermediary servers (such as Domain Name System (DNS) resolver, Network Time Protocol server, and Simple Service Discovery Protocol services) to magnify attack traffic without the need to compromise a large number of devices directly [3, 4].
To combat these diverse and evolving DDoS attack methods, a plethora of machine learning (ML) models have been developed to detect and mitigate DDoS attacks for specific network types (for IoT network traffic [5], software-defined networking traffic [6, 7], and tradition network traffic, c.f., [8]). Although these approaches show good results in detecting DDoS attacks on specific network types, they struggle in handling complex, heterogeneous environments like multi-environment (M-En) networks [9, 10].
M-En networks, which integrate multiple network types under a single environment, present unique challenges for DDoS detection due to multiple challenges. First, there is a lack of benchmark datasets to train models effectively. Second, the traffic patterns vary widely across IoT and traditional networks, each with unique packet characteristics; as a result, a model trained on one network type, like IoT, may not effectively identify patterns in traditional traffic. Another key challenge is the bias toward specific network traffic; if one traffic dominates, the model may overfit this traffic type, reducing performance across other networks. These challenges make it difficult to create a unified model that accurately generalizes across all network environments [11].
@@ I Introduction
Several studies explore M-En network security and address these challenges through various approaches. For instance, Rustam et al. [10] propose a fully automated malicious traffic detection system (MTDS) that addresses the lack of M-En datasets by generating M-En traffic through a combined IoT and traditional network dataset (IoTID-20 [12] and UNWNB-15 [13]). They tackle traffic diversity by deploying a Moth Flame Optimizer to find optimal weights of an ML algorithm for both network traffic types. Similarly, another study by Rustam et al. [14] employs an optimization approach to create M-En traffic, including IoT and traditional network traffic. They propose a self-diverse ensemble model by combining three variations of random forests optimized with a particle swarm optimizer. Most current studies create M-En traffic by merging existing datasets, which may not capture real-time M-En network patterns, and combining features can result in the loss of meaningful traffic patterns. Furthermore, no study currently addresses continuous learning, which could lead to performance degradation of these models over time.
This study addresses the challenges in securing M-En networks against DDoS attacks in two main contributions. In the first contribution, we introduce a comprehensive approach to M-En dataset collection that leverages the DDoShield-IoT testbed [5], which integrates NS-3 (a discrete-event network simulator) with Docker containers. This testbed allows us to generate real-world M-En traffic by running actual binaries in a simulated network, producing both benign traditional traffic (FTP, HTTP, and RTMP-based video) and malicious IoT traffic derived from Mirai malware binaries. To achieve a complete range of traffic types—namely benign IoT and malicious traditional traffic—we incorporate external PCAP datasets from the literature, ensuring our final dataset accurately reflects diverse, real-world network conditions.
@@
Following dataset construction, we uniformly process all collected packets through a consistent feature extraction pipeline, capturing both packet-level and time-based statistical features. This uniformity ensures that traffic from all sources—DDoShield-IoT and external PCAP datasets—can be directly compared and integrated. As a result, the dataset provides a robust foundation for training ML models capable of detecting DDoS attacks in M-En networks, with improved reliability and generalizability compared to prior approaches that relied solely on merged feature sets without preserving raw packet-level information.
In the second contribution, we develop and implement a continuous learning technique that combines ML and human (expert) intervention. Then, we deploy a real-time, continuous learning approach using a multi-level framework involving two models: M1 and M2. M1 is a lightweight model optimized for continuous learning with low computational costs and is initially trained on a smaller selected samples dataset to allow faster adaptability to new patterns. M2, in contrast, is a more robust and complex model with higher accuracy, trained on a larger, more comprehensive dataset. When a new packets arrives, M1 attempts a prediction. If it is unable to make a confident prediction, the data is passed to M2 for further analysis. Should M2 also fail to reach a definitive prediction, the data is then flagged to human experts for manual verification. Based on the expert’s feedback, M1 undergoes retraining to integrate the new insights, allowing it to enhance its predictive accuracy over time by incorporating this human-reviewed data into its model.
In summary, the key contributions of this work are as follows:
  • It proposes a real-time Multi-level Framework (Multi-LF) for DDoS attack detection in M-En networks.
@@
  • It introduces an M-En traffic dataset constructed by leveraging the DDoShield-IoT testbed and integrating external PCAP datasets. This dataset encompasses both IoT and traditional (non-IoT) traffic, including benign and malicious categories, thereby achieving a comprehensive and realistic representation of M-En network conditions.
  • A feature engineering approach is proposed to extract both general and statistical features by collecting packets within a specific time frame. Combining statistical features (e.g., Packet Counts, Destination Port Entropy) with general features creates a more linearly separable feature space, leading to substantial performance gains.
  • Multi-LF is tested and validated in real-time scenarios by deploying it in NS-3 simulations. This approach has demonstrated greater accuracy compared to state-of-the-art algorithms.
  • The performance of all models is evaluated based on accuracy, precision, recall, F1 score, model size, memory usage, CPU utilization, and prediction time.
@@
The rest of this paper is organized as follows: Section II presents the related work in the problem domain. Section III describes the proposed methodology. Section IV discusses the results, and Section VI provides the conclusion and future work.
II Related Work
@@
This section discusses the related works on malicious traffic detection for traditional networks, IoT networks, and multi-environment (M-En) networks. Furthermore, we explore the literature on continuous learning and identify gaps addressed by our study.
II-A Malicious Traffic Detection using Machine Learning
@@
Malicious traffic, particularly Distributed Denial-of-Service (DDoS) attacks, poses significant threats to network security. Many researchers have worked on efficiently tackling DDoS. For example, Anley et al. [15] propose an innovative approach that utilizes Convolutional Neural Networks (CNNs), adaptive architectures, and transfer learning techniques to detect malicious traffic. Their method demonstrates robust detection capabilities across various attack categories and is validated on publicly available datasets such as CIC-DDoS2019 [16], CSE-CIC-IDS2018 [17], UNSW-NB15 [13], and KDDCup’99 [18]. Their approach achieved accuracies of 93.62%, 99.92%, 99.84%, and 98.99% on each dataset, respectively. In a comparative study, Al-Eryani et al. [19] evaluate machine learning (ML) algorithms using the CIC-DoS2019 dataset [16], finding that Gradient Boosting (GB) and XGBoost achieve high accuracy (GB: 99.99%, XGBoost: 99.98%) with minimal false alarms.
@@
Persistent challenges in DDoS defense stem from evolving attack vectors and the increasing complexity of network environments. In response, Zhao et al. [1] have developed DFNet, an approach that integrates advanced ML models with packet scheduling algorithms in the network data plane. This approach effectively forwards 99.93% of victim-desired traffic during new DDoS attacks while incurring minimal overhead. Similarly, Singh et al. [20] contribute to the detection and mitigation of DDoS attacks in software-defined networking (SDN) environments. Their work also presents a new DDoS dataset with over 1.7 million entries and employs two detection methods: Snort [21] (an Intrusion Detection System) and eight different ML algorithms, including Ensemble Classifiers and a Hybrid Support vector machine-Random Forest (SVM-RF) classifier. The detection methods achieved 99.1% accuracy. In addition, the authors suggested two strategies to mitigate DDoS traffic: dropping illegitimate traffic and redirecting it.
II-B Malicious Traffic Detection in Traditional IP-based Network
@@
Traditional IP-based networks—networks that primarily handle non-IoT traffic using the Internet Protocol, distinguishing them from networks that may employ protocols like MQTT for IoT devices—have been extensively explored, resulting in numerous benchmark datasets and efficient security approaches. In recent literature, several researchers have proposed methods to protect these networks from malicious actors. Talukder et al. [22] present an ML-based network intrusion detection model that integrates Random Oversampling (RO) to counter data imbalance, Stacking Feature Embedding derived from clustering results, and Principal Component Analysis (PCA) for dimension reduction. Their model achieves exceptional accuracy rates: 99.59% and 99.95% with RF and Extra Trees (ET) models on the UNSW-NB15 dataset [13], 99.99% on the CIC-IDS-2017 dataset [23], and 99.94% on the CIC-IDS-2018 dataset [17] with DT and RF models, respectively.
@@
Expanding the focus on network security, Casanova et al. [24] concentrate on transforming the CIRACIC-DoHBrw-2020 time-series dataset [24] for training deep learning models in network intrusion detection. Their approach includes a two-layer network classification strategy, distinguishing DNS over HTTPS (DoH) from non-DoH traffic and further classifying DoH traffic into benign and malicious categories using a subset of 26 features and various types of Recurrent Neural Networks (RNNs), such as Long Short-Term Memory (LSTM), Bidirectional LSTM, Gated Recurrent Unit (GRU), and Deep RNN. Bi-LSTM outperforms all other methods with 99% accuracy, while GRU is the second-best performer. Hema et al. [25] propose a novel feature selection metric, CorrAUC, and develop a new feature selection algorithm using a wrapper technique. Their approach enhances traffic flow classification accuracy, evaluated using the NSL-KDD dataset [18] with three different ML algorithms such as RF, LR, and KNN, achieving 99%, 82%, and 98%, respectively. Babayigit et al. [26] propose the Queried Adaptive Random Forests (QARF) method—an online active learning-based approach that combines adaptive RF with an adaptive margin sampling strategy. This method queries a small number of instances from unlabeled traffic streams to obtain training data. Experimental evaluations using the NSL-KDD dataset [18] demonstrate that QARF achieves 98.20% accuracy.
II-C Malicious Traffic Detection in IoT Networks

### v2

# Multi-LF: A Continuous Learning Framework for Real-Time Malicious Traffic Detection in Multi-Environment Networks

@@ Abstract
Multi-environment (M-En) networks integrate diverse traffic sources, including Internet of Things (IoT) and traditional computing systems, creating complex and evolving conditions for malicious traffic detection. Existing machine learning (ML)-based approaches, typically trained on static single-domain datasets, often fail to generalize across heterogeneous network environments. To address this gap, we develop a realistic Docker–NS3-based testbed that emulates both IoT and traditional traffic conditions, enabling the generation and capture of live, labeled network flows. The resulting M-En Dataset combines this traffic with curated public PCAP traces to provide comprehensive coverage of benign and malicious behaviors. Building on this foundation, we propose Multi-LF, a real-time continuous learning framework that combines a lightweight model (M1) for rapid detection with a deeper model (M2) for high-confidence refinement and adaptation. A confidence-based coordination mechanism enhances efficiency without compromising accuracy, while weight interpolation mitigates catastrophic forgetting during continuous updates. Features extracted at 1-second intervals capture fine-grained temporal patterns, enabling early recognition of evolving attack behaviors. Implemented and evaluated within the Docker–NS3 testbed on live traffic, Multi-LF achieves an accuracy of 0.999 while requiring human intervention for only 0.0026% of packets, demonstrating its effectiveness and practicality for real-time malicious traffic detection in heterogeneous network environments.
keywords:
@@
Malicious Traffic Detection, Dataset Collection, M-En Networks, Continuous Learning, Zero-Day Attacks
[…]
[label1]organization=School of Computer Science, University College Dublin, addressline=, city=Dublin, postcode=D06WK27, country=Ireland
[…]
[label2]organization=Department of Computer Systems Technology, North Carolina A&T State University, city=North Carolina, country=USA
1 Introduction
Modern networks face a broad spectrum of malicious activities that now extend far beyond volumetric denial of service floods to include data exfiltration, protocol abuse, and other adaptive threats [10380450]. These attacks originate from an increasingly diverse set of hosts, including cloud servers, enterprise endpoints, and resource-constrained Internet of Things (IoT) devices, all interconnected in complex multi-environment (M-En) networks [NAZIR2023101820]. The resulting variety in traffic patterns, timing characteristics, and device capabilities undermines traditional malicious detection methods that depend on static signatures or models tuned to a single operating context. The challenge is compounded by (i) the heterogeneity between IoT and traditional traffic, (ii) the absence of a realistic M-En dataset, (iii) the class imbalance that causes models to overfit the environment that dominates the training data  [rustam2024ai], and (iv) the need for continuous monitoring with human expertise during online learning [rustam2025adaptive]. Effective defense, therefore, requires a framework that can ingest live traffic from multiple sources (e.g., traditional systems and IoT devices) and continuously learn to adapt in real time to unseen traffic, all while imposing minimal overheads [ismail2021review, 8299485].
@@ 1 Introduction
Researchers have addressed this emerging threat landscape by developing numerous machine learning (ML) models that target specific deployment contexts [MAHBOUBI2024104004]. Dedicated models have been trained for IoT traffic [devivo2024ddoshield], for flows within software-defined networking fabrics [9527066, BHAYO2023106432], and for conventional enterprise and backbone traffic (cf., [najafimehr2023ddos]). These models achieve high accuracy within the intended domain; however, their performance degrades when encountering traffic that mixes protocols, spans diverse device classes, or carries attack patterns it has never seen during training. Similarly, several studies have explored M-En network security and addressed these challenges through various approaches. For instance, Rustam et al. [rustam2024famtds] propose a fully automated malicious traffic detection system that addresses the lack of M-En datasets by generating M-En traffic through a combined IoT and traditional network dataset (IoTID-20 [IoTID20] and UNWNB-15 [7348942]). They tackle traffic diversity by deploying a moth flame optimizer (MFO) to find optimal weights of an ML algorithm for different types of network traffic. Similarly, another study by Rustam et al. [rustam2024malicious] employs an optimization approach to generate M-En traffic, which includes both IoT and traditional network traffic. They propose a self-diverse ensemble model by combining three variations of random forests optimized with a particle swarm optimizer (PSO). Nonetheless, most current studies create M-En traffic by merging existing datasets and then training and testing their models on the resulting dataset, which may not capture real-time M-En network patterns. In addition, combining the feature sets of different networks (since each dataset provides distinct features) can lead to the loss of meaningful traffic patterns. Lastly, none of these studies employs continuous learning with proper human oversight, causing their models’ performance to degrade over time.
@@
This paper addresses the challenges in securing M-En networks against malicious attacks in two main contributions. In the first contribution, we introduce a comprehensive approach to M-En dataset collection that utilizes the DDoShield-IoT testbed [devivo2024ddoshield], which leverages the integration of NS-3 [riley2010ns] (a network simulator) with Docker [merkel2014docker] (a container-based virtualization technology) to host real-world binaries that communicate over a simulated network. This testbed enables us to generate realistic M-En traffic by running actual binaries in a simulated network, producing benign and malicious traffic from the traditional and IoT networks, respectively. To achieve a complete range of traffic types (i.e., to incorporate the missing benign IoT and malicious traditional traffic), we utilize external PCAP datasets from the literature, ensuring our final dataset accurately reflects diverse and realistic M-En network conditions. Following our M-En dataset construction, we uniformly process all dataset packets through a consistent feature extraction pipeline, capturing both packet-level and time-based statistical features. This uniformity ensures that traffic from all sources can be directly compared and integrated. As a result, our M-En dataset provides a robust foundation for training ML models capable of detecting malicious attacks in M-En networks, offering improved reliability and generalizability compared to prior approaches that relied solely on merged feature sets without preserving raw packet-level information.
In the second contribution, we design and implement a continuous learning, real-time Multi-Level Framework (Multi-LF). Multi-LF combines two ML models, M1 and M2, and human (expert) intervention. Specifically, M1 is a lightweight model that is initially trained on a smaller, selected sample dataset. Furthermore, M1 is optimized for continuous learning with low computational costs, enabling adaptability to new traffic patterns and a faster detection rate with minimal resource requirements. M2, in contrast, is a more robust and complex model with higher accuracy, trained on a larger, more comprehensive dataset. In essence, Multi-LF operates as follows: when a new packet arrives, M1 attempts the first prediction; if its prediction confidence falls below a threshold, Multi-LF forwards the information about this packet to M2 for a second prediction attempt. Similarly, if M2 fails to reach a definitive prediction, Multi-LF escalate it for manual inspection by human experts. Then, Multi-LF updates M1 with the expert’s label using a weight interpolation method where 75% of M1’s existing weights remain intact and 25% come from a quick gradient update on the newly labeled sample. The updates to M1 introduce new knowledge while mitigating overfitting and catastrophic forgetting, which keeps M1 fast, adaptive, and accurate as traffic evolves.
We implement and deploy Multi-LF in DDoShield-IoT to assess its performance against live traffic. Using this deployment, we achieve an accuracy of 0.999. Across fifteen streaming batches covering more than 115,795 packets, only three instances required human intervention, resulting in just 0.0026% human involvement. Even with a high system depth, the latency rate remains unaffected. In addition, we measure resource usage and find that the lightweight M1 model maintains a throughput of 66,130 packets per second. In summary, the key contributions of this work are as follows.
  • We propose Multi-LF, a continuous learning, real-time framework that combines ML techniques with expert feedback for malicious attack detection in M-En networks. Code and dataset for reproducibility are available on GitHub [^1] https://github.com/furqanrustam/MultiLF.
@@
  • We construct an M-En traffic dataset by leveraging DDoShield-IoT and integrating external PCAP datasets. This dataset encompasses both IoT and traditional (non-IoT) traffic, including benign and malicious categories.
  • We propose transformer-based lightweight and complex model architectures, trained on a selective subset and a large corpus, respectively. Under Multi-LF, these models achieve significant performance with high throughput while imposing minimal burden on human workload.
  • We test and validate Multi-LF in real-time scenarios by hosting its implementation in DDoShield-IoT. This process demonstrates that Multi-LF results in higher accuracy compared to recent approaches.
  • We collect a wide range of performance metrics (e.g., accuracy, precision, throughput, CO2 emissions) and resource measurements (e.g., memory usage and CPU utilization) under real operating conditions.
@@
The rest of this paper is organized as follows. Section 2 presents the related work in the problem domain. Section 3 describes the proposed methodology. Section 4 discusses the evaluation and results, and Section 6 provides the conclusion and future work.
2 Related Work
@@
In this section, we discuss recent literature on malicious traffic detection for traditional, IoT, and M-En networks. Furthermore, we review the literature on continuous learning and identify gaps that our study addresses.
2.1 Malicious Traffic Detection using Machine Learning
@@
Malicious traffic, e.g., Distributed Denial of Service (DDoS) attacks, poses significant threats to network security. Many researchers have worked on efficiently mitigating these attacks. For example, Anley et al. [anley2024robust] propose an innovative approach that utilizes Convolutional Neural Networks (CNNs), adaptive architectures, and transfer learning techniques to detect malicious traffic. Their method demonstrates robust detection capabilities across various attack categories and is validated on publicly available datasets such as CIC-DDoS2019 [8888419], CSE-CIC-IDS2018 [Sharafaldin2018TowardGA], UNSW-NB15 [7348942], and KDDCup’99 [tavallaee2009detailed]. Their approach achieved accuracies of 93.62%, 99.92%, 99.84%, and 98.99% on each dataset, respectively. In a comparative study, Al-Eryani et al. [10485168] evaluate machine learning (ML) algorithms using the CICDoS2019 dataset [8888419], finding that Gradient Boosting (GB) and XGBoost achieve high accuracy (GB: 99.99%, XGBoost: 99.98%) with minimal false alarms.
Similarly, the study [10948513] proposes a graph transformer-based autoencoder (GTAE-IDS), an unsupervised packet-based graph intrusion detection framework designed for early and accurate anomaly detection. The system captures traffic and constructs sequential packet graphs, learning benign behavior using a graph autoencoder with transformer-based encoding to model contextual packet relationships. Unlike prior flow-based or supervised approaches, GTAE-IDS operates in near real time, requires no labeled data, and is resilient to unseen attacks. Experiments on CICIDS2017, CICIDS2018, and ACIIoT2023 show >>98% accuracy, even when observing only the first few packets of each communication. Another study [11199889] introduces a scalable SDN-CFN network model integrated with a GNN–based anomaly detection framework (GNN-NAD). The proposed system combines software-defined networking (SDN) and compute-first networking (CFN) to enable intelligent, real-time traffic analysis. GNN-NAD fuses static, vulnerability-aware attack graphs with dynamic traffic features, providing a unified and context-rich view of network security. At its core, a GNN model (GSAGE) performs graph representation learning, and an RF classifier conducts final detection. The hybrid design (GSAGE + RF) achieves a significant accuracy of 99.69, even with limited samples from the CICIDS2017 dataset.
@@
Persistent challenges in network defense stem from evolving attack vectors and the increasing complexity of network environments. In response, Zhao et al. [10380450] have developed DFNet, an approach that integrates advanced ML models with packet scheduling algorithms in the network data plane. This approach effectively forwards 99.93% of victim-desired traffic during new attacks while incurring minimal overhead. Similarly, Singh et al. [singh2024novel] contribute to the detection and mitigation of different attacks in software-defined networking (SDN) environments. Their work also presents a new dataset with over 1.7 million entries and employs two detection methods: Snort [waleed2022open] (an Intrusion Detection System) and eight different ML algorithms, including Ensemble Classifiers and a Hybrid Support Vector Machine-Random Forest (SVM-RF) classifier. The detection methods achieved 99.1% accuracy. In addition, the authors suggested two strategies to mitigate malicious traffic: dropping illegitimate traffic and redirecting it.
2.2 Malicious Traffic Detection in Traditional IP-based Network
@@
Traditional IP-based networks (networks that primarily handle non-IoT traffic using the Internet Protocol, distinguishing them from networks that may employ protocols like MQTT for IoT devices) have been extensively explored, resulting in numerous benchmark datasets and efficient security approaches. In recent literature, several researchers have proposed methods to protect these networks from malicious actors. Talukder et al. [talukder2024machine] present an ML-based network intrusion detection model that integrates Random Oversampling (RO) to counter data imbalance, Stacking Feature Embedding derived from clustering results, and Principal Component Analysis (PCA) for dimension reduction. Their model achieves exceptional accuracy rates: 99.59% and 99.95% with RF and Extra Trees (ET) models on the UNSW-NB15 dataset [7348942], 99.99% on the CIC-IDS-2017 dataset [sharafaldin2019detailed], and 99.94% on the CIC-IDS-2018 dataset [Sharafaldin2018TowardGA] with DT and RF models, respectively.
@@
Expanding the focus on network security, Casanova et al. [24] concentrate on transforming the CIRACIC-DoHBrw-2020 time-series dataset [9878102] for training deep learning models in network intrusion detection. Their approach includes a two-layer network classification strategy, distinguishing DNS over HTTPS (DoH) from non-DoH traffic and further classifying DoH traffic into benign and malicious categories using a subset of 26 features and various types of Recurrent Neural Networks (RNNs), such as Long Short-Term Memory (LSTM), Bidirectional LSTM, Gated Recurrent Unit (GRU), and Deep RNN. Bi-LSTM outperforms all other methods with 99% accuracy, while GRU is the second-best performer. Hema et al. [hema2023malicious] propose a novel feature selection metric, CorrAUC, and develop a new feature selection algorithm using a wrapper technique. Their approach enhances traffic flow classification accuracy, evaluated using the NSL-KDD dataset [tavallaee2009detailed] with three different ML algorithms, such as RF, LR, and KNN, achieving 99%, 82%, and 98%, respectively. Babayigit et al. [10543240] propose the Queried Adaptive Random Forests (QARF) method, an online active learning-based approach that combines adaptive RF with an adaptive margin sampling strategy. This method queries a small number of instances from unlabeled traffic streams to obtain training data. Experimental evaluations using the NSL-KDD dataset [tavallaee2009detailed] demonstrate that QARF achieves 98.20% accuracy.
2.3 Malicious Traffic Detection in IoT Networks